### v2

# Connectedness of the Gromov boundary of fine curve graphs

Abstract
In this paper, we study the topological properties of the Gromov boundary of the fine curve graph of an orientable finite-type surface of genus at least 2. This graph consisting of topological curves has much richer dynamics than the classical curve graph. Using the techniques introduced by Wright [Wri23], we show that this boundary is (path) connected and that the spheres in non-separating fine curve graph are connected.
Keywords: fine curve graph, Gromov boundary, connectedness.
2020 Mathematics subject classification: 57K20, 53C23.
1 Introduction
Let $S$ be an orientable finite-type surface of genus $g\geq 2$. In order to study the group of diffeomorphisms of surface $G$, Bowden, Hensel and Webb design for $S$ its fine curve graph $\mathcal{C}^{\dagger}(S)$ [BHW22], where the vertices are smooth curves instead of their isotopy classes, and the edges are defined by disjointness.
Indeed, the fine curve graph $\mathcal{C}^{\dagger}(S)$ shares a lot of similarity with the curve graph of $S$ and can be considered as an analogue of the curve graph for the mapping class group: the graph is Gromov hyperbolic [BHW22] and the group of homeomorphisms of surface $S$ is isomorphic to the automorphism group of the topological version of the fine curve graph [Lon+21].
But compared to the classical curve graph, the fine curve graph $\mathcal{C}^{\dagger}(S)$ is much more complicated and the dynamics of homeomorphism group on $\mathcal{C}^{\dagger}(S)$ is much richer than the mapping class group on the curve graph: $\mathcal{C}^{\dagger}(S)$ is a locally uncountable; there exist homeomorphisms on the surface $S$ that induce parabolic isometries on $\mathcal{C}^{\dagger}(S)$ and by consequences there are points on the Gromov boundary that are not represented by ending laminations [BHW22, Bow+22], whereas it is the case for curve graphs [Kla22].
The study of geometric or topological interpretation of points at infinity in the fine curve graph is still an ongoing project. Efforts are made in understanding the stabilisers of these points in terms of (homological) rotation sets [Bow+22, GM23a, GM23].
In this paper, we are more interested in the topological properties of the Gromov boundary of fine curve graphs. This paper is written as one of the initiatives in this direction.
For the curve graph, Peter Storm asked an interesting question about the (path)-connectedness of its the Gromov boundary, which is recorded in [KL08, Question 10]. This question is answered positively by [Gab09, LS09, LMS11]. Recently, [Wri23] provides a new proof and proves that the boundary of curve graph is in addition linearly connected (see Definition 2.10).
Let $\operatorname{\mathcal{NC}}^{\dagger}(S)$ be the subgraph of $\mathcal{C}^{\dagger}(S)$ consisting of non-separating curves.
In this paper, we conclude the following results for the topology of Gromov boundary of the fine curve graph:
Theorem 1.1.
Let $S$ be an orientable finite-type surface of genus $g\geq 2$. Then the Gromov boundary $\partial\mathcal{C}^{\dagger}(S)$ of the fine curve graph is path connected.
Moreover, for any $o\in\operatorname{\mathcal{NC}}^{\dagger}(S)$ and any small enough $\varepsilon>0$, if $b\in(1,1+\varepsilon)$ and $\rho_{o,b}$ is the visual metric defined on the Gromov boundary, then it is linearly connected.
We remark that Theorem 1.1 cannot be deduced from the path-connectedness of the Gromov boundary of curve graphs, since there are boundary points in the fine curve graph that are cannot be represented by ending laminations and that path-connectedness is usually not preserved under taking closures.
We say that a subgraph $G$ of a graph $\Gamma$ is connected if for any two points $a,b\in G$, there exists a finite sequence $a=v_{0},\ v_{1},\ \dots,\ v_{n}=b$ such that $v_{i}$ is connected to $v_{i+1}$ by an edge for all $10\leq i\leq n-$ 1 and $v_{i}\in G$ for all $0\leq i\leq n$.
Theorem 1.2.
Let $S$ be an orientable finite-type surface of genus $g\geq 2$. For any integer $r>0$ and any $o\in\operatorname{\mathcal{NC}}^{\dagger}(S)$, the $r$-sphere $\{\alpha\in\operatorname{\mathcal{NC}}^{\dagger}(S)\colon d^{\dagger}(\alpha,o)=r\}$ is connected.
The main difficulties of dealing with the fine curve graph is the non-transversality between topological curves and the lack of analogue in the setting of fine curve graphs of tight geodesics as in the curve graphs. In order to deduce the results, we first generalise the criterion introduced by Wright in [Wri23] to any hyperbolic graphs with no “dead ends” (see Proposition 2.11). In order to apply Proposition 2.11, it suffices to check that the assumptions are satisfied by a subgraph quasi-isometric to $\mathcal{C}^{\dagger}(S)$. In principle, under the non-separating and transverse settings, as long as there are only finitely many topological curves involved, the fine curve graph looks like the curve graph of a punctured subsurface (see Proposition 3.5), which allows us to make use of [Wri23, Proposition 5.4]. The technical part of the proof lies in passing non-transverse pairs of curves into transverse ones. At the end of this paper, we also gave several interesting observations on the group action of $\operatorname{Homeo}(S)$ on the Gromov bordification of $\mathcal{C}^{\dagger}(S)$.
Acknowledgement
The first author wishes to thank the hospitality of Guangxi Center for Mathematical Research for a visit, during which this work is partially carried out. The second author also thanks the hospitality of Tsinghua University and Sun Yat-sen University. The first author is supported by CSC-202108170018. The second author is supported by the National Natural Science Foundation of China (No. 12001122, 12271533, 12361014) and the special grants for Guangxi Ba Gui Scholars. The authors are grateful to Yongsheng Jia for informing discussion and to Alex Wright for explaining to us some detail in his work. The authors also thank Bruno Duchesne and Pierre-Antoine Guihéneuf for reviewing the early drafts of this paper.
2 Gromov hyperbolic space and bordification
The notion of $\delta$-hyperbolic space is originally introduced in [Gro87]. Classical references are [BH13] and [DSU17], while the latter treats more general cases. Here we cover a very limited part of this topic to make clarified some folklore results and introduce the tools that we will later need. For more details, we refer the reader, if interested, to the classical references above.
In the sequel, we will write $O(\delta)$ for an additive error that is at most a multiple of $\delta>0$, i.e. $f=g+O(\delta)$ if and only if there exists $K>0$ such that $\|f-g\|_{\infty}<K\delta$.
Let $(X,d_{X})$ be a metric space. Let $o\in X$ be an arbitrary base point. Define the function $\langle\cdot,\cdot\rangle_{o}\colon X\times X\to[0,\infty)$ by
$\langle x,y\rangle_{o}\coloneqq\frac{1}{2}\big{(}d_{X}(x,o)+d_{X}(y,o)-d_{X}(x,y)\big{)}$ (1)
for every $x,y\in X$. The quantity $\langle x,y\rangle_{o}$ in (1) is called the Gromov product of $x$ and $y$ based at $o$, or simply the Gromov product of $x$ and $y$.
Definition 2.1 (Gromov hyperbolic space).
A metric space $(X,d_{X})$ is Gromov hyperbolic, or $\delta$-hyperbolic, if there exists some $\delta\geq 0$ such that
$\langle x,y\rangle_{o}\geq\min\left\{\langle x,z\rangle_{o},\langle z,y\rangle_{o}\right\}-\delta$ (2)
for all $x,y,z,o\in X$.
When the metric space is geodesic, i.e. a space where every pair of points can be connected by a path isometric to a closed real interval, $\delta$-hyperbolicity is equivalent to the property that every geodesic triangle is $\delta^{\prime}$-slim, for some uniform $\delta^{\prime}\geq 0$. This condition is attributed to Rips by Gromov in [Gro87] and is thus also often called Rips condition, see [DSU17, §4.3] for the exact definition.
A sequence $(x_{n})_{n\geq 0}$ in $\delta$-hyperbolic space is Cauchy-Gromov if $\langle x_{n},x_{m}\rangle_{o}\to\infty$ as $n,m\to\infty$. We remark that being Cauchy-Gromov of a sequence does not depend on the choice of the base point $o\in X$. A Cauchy-Gromov sequence $(x_{n})_{n\geq 0}$ is equivalent to another Cauchy-Gromov sequence $(y_{m})_{m\geq 0}$, if $\langle x_{n},y_{m}\rangle_{o}\to\infty$ as $n,m\to\infty$.
Definition 2.2 (Gromov boundary).
Let $(X,d_{X})$ be a Gromov hyperbolic space. Then we define its Gromov boundary by the equivalent classes of Cauchy-Gromov sequences in $X$ and we denote it by $\partial X$. We say that a sequence $(x_{n})_{n\geq 0}$ converges to a boundary point $\xi\in\partial X$ if it is a representative of $\xi$. The disjoint union $\mathrm{bord}X\coloneqq X\cup\partial X$ is called the Gromov bordification of $X$.
The Gromov product can be extended to the boundary in the following way. For all $x\in X$ and $\xi\in\partial X$, we define
$\langle x,\xi\rangle_{o}\coloneqq\inf\left\{\liminf_{n\to\infty}\langle x,x_{n}\rangle_{o}\colon(x_{n})\in\xi\right\},$
and also by setting for all $\eta,\xi\in\partial X$
$\langle\eta,\xi\rangle_{o}\coloneqq\inf\left\{\liminf_{n,m\to\infty}\langle y_{m},x_{n}\rangle_{o}\colon(x_{n})\in\xi\mathrm{\ and\ }(y_{m})\in\eta\right\}.$
In particular, note that $\langle\xi,\xi\rangle_{o}=\infty$ for all $\xi\in\partial X$. Moreover, we have the estimation
$\left|\lim_{n,m\to\infty}\langle x_{n},y_{m}\rangle_{o}-\langle\xi,\eta\rangle_{o}\right|=O(\delta)$ (3)
for any Cauchy-Gromov sequences $(x_{n})\in\xi$ and $(y_{n})\in\eta$ and any $o\in X$ [DSU17, Lemma 3.4.7].
With Gromov product, the Gromov boundary can be endowed with the uniform structure generated by the entourages in the form of
$\left\{(\xi,\eta)\in\partial X\times\partial X\colon\langle\xi,\eta\rangle_{o}>R>0\right\},$
which further makes it into a topological space. Another way to understand the open sets in $\partial X$ is to consider the basis consisting of shadows, i.e. subsets given by
$\left\{\xi\in\partial X\colon\langle x,\xi\rangle_{o}>R>0\right\}.$ (4)
This fact can be easily verified by (2) and (3) under the settings of Gromov hyperbolicity.
Recall that a map $f\colon(X,d_{X})\to(Y,d_{Y})$ between two metric spaces is a $(\lambda,k)$-quasi-isometric embedding if there exists $\lambda\geq 1$ and $k>0$ such that
$\frac{1}{\lambda}d_{X}(x,y)-k\leq d_{Y}\big{(}f(x),f(y)\big{)}\leq\lambda d_{X}(x,y)+k$
for every $x,y\in X$. If in addition, the map $f$ is essentially surjective, i.e. there exists $C>0$ such that $d_{Y}\big{(}y,f(X)\big{)}<C<\infty$ for all $y\in Y$, then $f$ is a quasi-isometry and we say that $X$ and $Y$ are quasi-isometric. We remark that being quasi-isometric is an equivalent relation. A quasi-geodesic in a metric space $X$ is a quasi-isometric embedding of a real interval into $X$.
The following result is pivotal in large scale geometry à la Gromov and is well-known in proper cases. But we remark that it actually holds for non-proper spaces [Has22, Väi05]. For convenience reason, we state it here below. There are several different proofs from different approaches and the one we present here might have already been known to [Cap+15]:
Proposition 2.3.
Let $(X,d_{X})$ be a $\delta$-hyperbolic geodesic space and $o\in X$ be an arbitrary base point. Then $\partial X$ is in bijection with the equivalent classes of quasi-geodesic rays issued from $o$ and two quasi-geodesic rays are equivalent if they have bounded Hausdorff distance. Moreover, a quasi-isometric embedding $f\colon X\to Y$ induces a topological embedding $f_{\partial}\colon\partial X\to\partial Y$ and $f_{\partial}$ is a homeomorphism if $f$ is essentially surjective.
Sketch of Proof.
First, fix a base point $o\in X$. We remark that for every boundary point $\eta\in\partial X$, we can connect $o$ to $\xi$ by a continuous $(1,N\delta)$-quasi-geodesic for some large $N>0$. Let $\omega$ be a non-principal ultrafilter over $\mathbb{N}$. Take the ultraproduct $u(X)$ of $X$ along $\omega$ in the following way: $u(X)$ is the space consisting of all equivalent bounded sequences in $X$ and two sequences $(x_{n})$ and $(y_{n})$ are equivalent if the $\omega$-limit of $d_{X}(x_{n},y_{n})$ tends to 00. The space $u(X)$ is equipped with the limit metric $d_{\omega}([x],[y])\coloneqq\lim_{n\to\omega}d_{X}(x_{n},y_{n})$. It is not hard to see that with this metric, $u(X)$ is geodesic and $\delta$-hyperbolic, where $X\hookrightarrow u(X)$ is embedded isometrically. Moreover, one can verify that some geodesic segments with endpoints on a quasi-geodesic ray or a bi-infinite quasi-geodesic line converge to a geodesic ray or a geodesic line. This fact plays the rôle of Arzelà-Ascoli theorem as in the classical proper cases. Now we can conclude the results, including Morse Lemma for infinite quasi-geodesics, as in the classical cases. ∎
Remark 2.4.
In [Gro87, §7], a Gromov hyperbolic space $X$ is ultracomplete if for any two points $x,y\in\mathrm{bord}X$, there exists a geodesic connecting $x$ to $y$. The proof above actually shows that $u(X)$ is ultracomplete. If a Gromov hyperbolic geodesic space is proper, then it is ultracomplete; but the converse is not true: the curve graph of a compact hyperbolic surface is ultracomplete but not proper [Min10, Lemma 5.14].
Another powerful tool to study the Gromov boundary is the following notion of visual metric that is carefully treated in [DSU17, §3.6] and [Väi05] under general settings:
Theorem 2.5 (Väsäilä).
Let $(X,d_{X})$ be a Gromov hyperbolic space. Then $\partial X$ is completely metrizable. For any based point $o\in X$, the topology is compatible to the visual metric $\rho_{o,b}$ from $o\in X$ with parameter $b\in(1,1+\varepsilon)$ for some small enough $\varepsilon>0$. Moreover, we have the estimation
$b^{-\langle\xi,\eta\rangle_{o}}/4\leq\rho_{o,b}(\xi,\eta)\leq b^{-\langle\xi,\eta\rangle_{o}}$ (5)
for every $\xi,\eta\in\partial X$.
Let $X$ be a $\delta$-hyperbolic geodesic space and let $\gamma\subset X$ be a geodesic. We define for a point $x\in X$ its nearest points projection $\pi_{\gamma}(x)$ by the points $y\in\gamma$ satisfying $d_{X}(x,y)=d_{X}(x,\gamma)$. We remark that $\pi_{\gamma}(x)$ is not empty in a geodesic space: it consists of the minima of the continuous proper function $d_{X}(x,\cdot)\colon\gamma\to[0,\infty)$.
The first result about this notion is referred as the reverse triangle inequality:
Proposition 2.6 ([MT18], Proposition 2.2).
Let $X$ be $\delta$-hyperbolic and $\gamma$ be a geodesic. For any $x\in X$ and any $p\in\pi_{\gamma}(x)$, we have
$d(x,y)=d(x,p)+d(p,y)+O(\delta)$ (6)
for all $y\in\gamma$.
Clearly from (6), the nearest point projection is coarsely well-defined: the diameter of $\pi_{\gamma}(X)$ is bounded uniformly by a multiple of $\delta$.
Using this inequality, we can deduce the following estimation, which is essentially a translation into nearest points projection language of the thinness of geodesic triangles in Gromov hyperbolic spaces:
Corollary 2.7.
Let $X$ be a geodesic $\delta$-hyperbolic space. Let $x,y,o\in X$ be three points and $\gamma\coloneqq[o,y]$ be a geodesic segment. Then for any $p\in\pi_{\gamma}(x)$,
$\langle x,y\rangle_{o}=d_{X}(o,p)+O(\delta).$ (7)
Proof.
Using (6) for $d_{X}(x,o)$ and $d_{X}(x,y)$ appeared in (1) will yield the result. ∎
Now we can prove the following lemma about the distance between two boundary points under a visual metric:
Lemma 2.8.
Let $X$ be a $\delta$-hyperbolic geodesic space. Fix a base point $o\in X$. Let $\rho_{o,b}$ be a visual metric from $o$ with parameter $b\in(1,1+\varepsilon)$ for sufficiently small $\varepsilon>0$. Then for every large enough $E\gg 1$, $\lambda\geq 1$ and $k\geq 0$, there exists a constant $C\coloneqq C(b,k,E,\delta)>0$ such that for every two distinct points $\xi,\eta\in\partial\Gamma$ and for any $\ell_{1},\ell_{2}$ continuous $(\lambda,k)$-quasi-geodesics connecting $o$ to $\xi$ and $\eta$ respectively, if
$r\coloneqq\sup\{t\geq 0\colon d_{X}(\ell_{1}(t),\ell_{2}(t))\leq E\},$
then
$b^{-\lambda r}/C\leq\rho_{o,b}(\xi,\eta)\leq Cb^{-r/\lambda}.$ (8)
Proof.
By continuity of $\ell_{1}$ and $\ell_{2}$, if $x\coloneqq\ell_{1}(r)$ and $y\coloneqq\ell_{2}(r)$, then $d_{X}(x,y)=E$. Let $\gamma\coloneqq[o,y]$ be a geodesic segment and $p\in\pi_{\gamma}(x)$. Applying the reverse triangle inequality (6) twice, we have
$d_{X}(p,y)=d_{X}(x,y)-d_{X}(x,p)+O(\delta)=E-d_{X}(x,o)+d_{X}(o,p)+O(\delta).$
Using the fact that $d_{X}(o,y)=d_{X}(o,p)+d_{X}(p,y)$, we can further deduce that
$d_{X}(o,p)=\frac{1}{2}\left(d_{X}(x,o)+d_{X}(y,o)-E\right)+O(\delta).$
Using the $(\lambda,k)$-quasi-geodesicity, we have the inequality
$\frac{r}{\lambda}-k-E/2+O(\delta)\leq d_{X}(o,p)\leq\lambda r+k-E/2+O(\delta).$ (9)
But by (3), when $E$ is sufficiently large, we have $\langle x,y\rangle_{o}=\langle\xi,\eta\rangle_{o}+O(\delta)$. Combining this fact with (7) and (9), we can deduce that
$\frac{r}{\lambda}-k-E/2+O(\delta)\leq\langle\xi,\eta\rangle_{o}\leq\lambda r+k-E/2+O(\delta).$
By setting $C\coloneqq 4b^{k+E/2+O(\delta)}$, we can deduce the desired inequality via applying the estimation (5). ∎
Remark 2.9.
As mentioned in the proof of Proposition 2.3, in a geodesic $\delta$-hyperbolic space, each boundary point $\xi\in\partial X$ can be connected to the base point $o$ by a continuous $(1,N\delta)$-quasi-geodesic. So the $\lambda$ from Lemma 2.8 can be taken as $1$.
At this moment, we introduce the following notion:
Definition 2.10 (Linear connectedness).
A metric space $(X,d_{X})$ is linearly connected if there is a constant $L>0$ such that for each pair $x,y\in X$ there is a compact connected set $K\subset X$ containing $x,y$ with diameter at most $Ld_{X}(x,y)$.
The application of Lemma 2.8 is the following proposition inspired by [MS20, Proof of Proposition 5.2] and is an improvement of [Wri23, Proposition 2.2].
For notations, let $S_{o}(r)$ be the sphere centred at $o$ of radius $r>0$ and $B_{o}(r)$ be the ball centred at $o$ of radius $r>0$.
Proposition 2.11.
Let $\Gamma$ be a Gromov hyperbolic simplicial graph and let $o\in\Gamma$ be a base point. Suppose that the following holds.
  1. (G1)
    Every vertex of $\Gamma$ is adjacent to point of $\Gamma$ that is $1$ farther from $o$.
If in addition there is some $D>0$ such that for all $r\geq 0$, the following conditions are satisfied:
  1. (G2)
    For every $z\in S_{o}(r)$ and $x,y\in S_{o}(r+1)\cap B_{z}(1)$ there exists a path
    $x=x_{0},\ x_{1},\ \dots,\ x_{\ell}=y$
    with
    $x_{i}\in\left(\Gamma-B_{o}(r)\right)\cap B_{x}(D)$
    for $0\leq i\leq\ell$.
  2. (G3)
    For every adjacent pair $x,y\in S_{o}(r)$ there exists a path
    $x=x_{0},\ x_{1},\ \dots,\ x_{\ell}=y$
    with
    $x_{i}\in\left(\Gamma-B_{o}(r)\right)\cap B_{x}(D)$
    for $0<i<\ell$.
Then equipped with the visual metric $\rho_{o,b}$ from $o$ with parameter $b\in(1,1+\varepsilon)$ for some small enough $\varepsilon>0$, the Gromov boundary $\partial\Gamma$ is linearly connected and path connected.
Proof.
The proof is a straightforward adaption of the proof of [Wri23, Proposition 2.2]. Let $\xi_{0},\xi_{1}\in\partial X$ be any two points on the Gromov boundary. Now connect $\xi_{0}$ and $\xi_{1}$ to $o$ respectively by continuous $(1,k)$-quasi-geodesic rays $\gamma_{0}$ and $\gamma_{1}$. By the same arguments in [Wri23, Proof of Proposition 2.2], (G1) assures the existence of a dense subset $I\subset[0,1]$ such that for all $t\in I$, there is a geodesic ray $\gamma_{t}\subset\Gamma$ connecting $o$ to a point $c(t)\in\partial\Gamma$. Since the family $\left(\gamma_{t}\right)_{t\in I\cup\{0,1\}}$ consists of continuous $(1,k)$-quasi-geodesic rays, thus by Lemma 2.8, we can deduce from (G2) and (G3) the same estimation as in [Wri23, Proof of Proposition 2.2], i.e. there exists $L>0$ independent of $\xi_{0}$ and $\xi_{1}$ such that the diameter $\mathrm{diam}\big{(}c(I)\big{)}<L\rho_{o,b}(\xi_{0},\xi_{1})$ and that $c\colon I\to\partial\Gamma$ is uniformly continuous. The desired path is then the closure of $c(I)$. ∎
Remark 2.12.
Proposition 2.11 cannot be obtained by applying [Wri23, Proposition 2.2].
3 Fine curve graph
Let $S$ be an orientable finite-type surface of genus $g\geq 2$. It can carry a hyperbolic metric. In a recent paper, Bowden, Hensel and Webb design for $S$ its fine curve graph [BHW22] as a combinatorial tool for studying $\operatorname{Homeo}(S)$, which is an analogue to the renown curve graph.
An essential simple closed smooth curve $\gamma$ is a proper $\mathcal{C}^{\infty}$-embedding of the circle $S^{1}\hookrightarrow S$ such that $\gamma$ is not homotopic to a singleton nor bounds a once-punctured disc or an annulus. If in addition $S-\gamma$ is connected, then the curve is non-separating.
Remark 3.1.
In what follows, all curves will be essential on the surface $S$, simple, closed, and smooth, unless otherwise mentioned.
The fine curve graph $\mathcal{C}^{\dagger}(S)$ is a graph where the vertices are essential simple closed differentiable curves on $S$ and two vertices are joined by an edge if they are disjoint. The non-separating fine curve graph $\operatorname{\mathcal{NC}}^{\dagger}(S)$ is the subgraph of $\mathcal{C}^{\dagger}(S)$ consisting of vertices as non-separating curves. If they are equipped with the simplicial metrics, then these two connected graphs are quasi-isometric (with the inclusion map being a $(1,1)$-quasi-isometry) and are simultaneously Gromov hyperbolic [BHW22]. Hence by Proposition 2.3, the Gromov boundary $\partial\operatorname{\mathcal{NC}}^{\dagger}(S)$ is homeomorphic to $\partial\mathcal{C}^{\dagger}(S)$.
Remark 3.2.
We remark that it is possible to construct the fine curve graphs by using non-differentiable curves, i.e. $\mathcal{C}^{0}$-embedding of $S^{1}\hookrightarrow S$, which turns out to be the topological version of the fine curve graphs. However, we remark that any topological curves can be isotopic to a differentiable curve that is disjoint from it: isotopically mollify the curve and then move it away. Hence there is a $1$-Lipschitz embedding of the fine curve graphs of differentiable curves into the fine curve graphs of non-differentiable curves and this embedding is $1$-dense, i.e. the two versions of fine curve graphs are quasi-isometric and have no difference in the view point of large-scale geometry.
In the sequel, we will denote by $d^{\dagger}$ the simplicial metric on $\operatorname{\mathcal{NC}}^{\dagger}(S)$. Also, we will also denoted by $\operatorname{\mathcal{NC}}(S)$ the simplicial graph where the vertices are isotopic classes of essential non-separating simple closed curves and two vertices are connected by an edge if they admit disjoint representatives. We remark that $\operatorname{\mathcal{NC}}(S)$ is also Gromov hyperbolic and is quasi-isometric to the usual curve graph [Ham14, §3].
Definition 3.3 (Transversality).
Let $\alpha,\beta$ be two simple closed curves. We say that they are transverse if exactly one of the following holds:
  • $\alpha\cap\beta=\emptyset$.
  • For every $p\in\alpha\cap\beta$, there is an open neighbourhood $N\subset S$ of $p$ and a diffeomorphism onto the unit disk $\varphi\colon N\to D\subset\mathbb{R}^{2}$ such that $\varphi(\alpha\cap N)=(-1,1)\times\{0\}$ and $\varphi(\beta\cap N)=\{0\}\times(-1,1)$.
Note that if two curves are transverse, then by compactness, they only intersect at finitely many points.
A bigon of two transverse curves $\alpha,\beta$ in a (punctured) surface is a connected component in the complementary $\alpha\cup\beta$ that is homeomorphic to a disk and bounds exactly one subarc of $\alpha$ and one subarc of $\beta$. Two transverse curves $\alpha,\beta$ are in minimal position if they have no bigon [FM11, Proposition 1.7].
One advantage of being smooth is that the curves can be taken as geodesic under some hyperbolic metric:
Proposition 3.4 (Proposition 10, Exposé 3, [FLP91]).
Let $S$ be an orientable finite-type surface of genus at least $2$. Let $\gamma$ be an essential simple closed smooth curve. Then there exists a hyperbolic metric on $S$ such that $\gamma$ is geodesic. If in addition $\gamma$ and $\gamma^{\prime}$ are two smooth curves in minimal positions, then there exists a hyperbolic metric on $S$ such that $\gamma$ and $\gamma^{\prime}$ are simultaneously geodesic.
An important metric geometric property of the fine curve graph is that it can be approximated via the curve graph of its punctured subsurfaces:
Proposition 3.5 (Lemma 3.4 & Lemma 3.6 [BHW22]).
Suppose that $\alpha,\beta\in\operatorname{\mathcal{NC}}^{\dagger}(S)$ are transverse, and that $\alpha$ and $\beta$ are in minimal position in $S-P$, where $P\subset S$ is finite and disjoint from $\alpha\cup\beta$. Then
$d_{\operatorname{\mathcal{NC}}(S-P)}([\alpha]_{S-P},[\beta]_{S-P})=d^{\dagger}(\alpha,\beta),$
where $[\alpha]_{S-P}$ and $[\beta]_{S-P}$ are respectively the isotopy classes of $\alpha$ and $\beta$ in $S-P$. Moreover, there is a geodesic $\alpha=\nu_{0},\ \nu_{1},\ \dots,\ \nu_{n}=\beta$ in $\operatorname{\mathcal{NC}}^{\dagger}(S)$ such that $\nu_{i}\cap P=\emptyset$ and $[\alpha]_{S-P}=[\nu_{0}]_{S-P},\ [\nu_{1}]_{S-P},\ \dots,\ [\nu_{n}]_{S-P}=[\beta]_{S-P}$ is also a geodesic in $\operatorname{\mathcal{NC}}(S-P)$.
Remark 3.6.
The finite set $P$ can be empty. In that case, it yields an isometric embedding of $\operatorname{\mathcal{NC}}(S-P)\hookrightarrow\operatorname{\mathcal{NC}}^{\dagger}(S)$. This can be realised by choosing a hyperbolic metric on $S$ and each isotopy class of simple closed curves is sent to its geodesic representative.
Now we will fix an arbitrary base point $o\in\operatorname{\mathcal{NC}}^{\dagger}(S)$. We write $\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$ the subgraph of $\operatorname{\mathcal{NC}}^{\dagger}(S)$ where all the vertices are transverse to $o$. It is a priori equipped with the subgraph distance but not the induced distance.
Lemma 3.7.
Let $\alpha\in\operatorname{\mathcal{NC}}^{\dagger}(S)$. Let $F\subset\operatorname{\mathcal{NC}}^{\dagger}(S)$ be a finite collection of pairwise transverse curves. Suppose that for every $\gamma\in F$, $\alpha\cap\gamma\neq\emptyset$. Then there exists a curve $\alpha^{\prime}\in\operatorname{\mathcal{NC}}^{\dagger}(S)$ isotopic to $\alpha$ that is disjoint from $\alpha$ and is transverse to all $\gamma\in F$.
Proof.
Let $A\subset S$ be a small enough annular neighbourhood of the curve $\gamma\in F$. Since $\alpha$ is a proper embedding of $S^{1}$, there are only finitely many connected components in $\alpha\cap A$. Hence it is possible to cut $A$ into finitely many pieces $A_{1},\ \dots,\ A_{n}$ so that for each $1\leq i\leq n$, the piece $A_{i}$ contains exactly one connected component of $\alpha\cap A$. Note that each $A_{i}$ is diffeomorphic to a $[-1,1]^{2}\subset\mathbb{R}^{2}$ with the image of $o\cap A_{i}$ being $[-1,1]\times\{0\}$. It is an elementary exercise to show that under this diffeomorphism, $\alpha$ is isotopic to an $\alpha^{\prime}$ that is transverse to $\gamma$ and disjoint from $\alpha$ inside of $[-1,1]^{2}$. Note that if $A$ is small enough, then $\alpha^{\prime}$ can be taken so that whenever $\alpha$ is transverse to some other finitely many curves, $\alpha^{\prime}$ remains transverse to them. Hence by induction, we can prove the desired result. ∎
We notice that the perturbation in Lemma 3.7 can be done within an arbitrarily small annular neighbourhood. All the isotopic perturbation in the sequel will be of the same flavour, although the situation might be slightly different.
Proposition 3.8.
The inclusion map $\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)\hookrightarrow\operatorname{\mathcal{NC}}^{\dagger}(S)$ is a $(1,2)$-quasi-isometry, hence $\partial\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)\simeq\partial\operatorname{\mathcal{NC}}^{\dagger}(S)$.
Proof.
For any $\alpha\in\operatorname{\mathcal{NC}}(S)$, if it is not transverse to $o$, then by Lemma 3.7, there is another $\alpha^{\prime}\in\operatorname{\mathcal{NC}}^{\dagger}(S)$ that is disjoint from $\alpha$ and is transverse to $o$. Hence $d^{\dagger}(\alpha,\alpha^{\prime})=d^{\dagger}\big{(}\alpha,\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)\big{)}=1$.
For any two vertices $\alpha,\beta\in\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$, if they are not transverse, then by applying Lemma 3.7, we can find a $\beta^{\prime}$ with $d^{\dagger}(\beta,\beta^{\prime})=1$ that is transverse to both $\alpha$ and $o$. Now we can find a geodesic $\alpha=\nu_{0},\ \nu_{1},\ \dots,\ \nu_{n}=\beta^{\prime}$ in $\operatorname{\mathcal{NC}}^{\dagger}(S)$ connecting $\alpha$ to $\beta^{\prime}$ with the property that there is a finite subset $P\subset S$ disjoint from $o$ such that $\nu_{i}\cap P=\emptyset$ for all $i$ and $[\alpha]_{S-P}=[\nu_{0}]_{S-P},\ [\nu_{1}]_{S-P},\ \dots,\ [\nu_{n}]_{S-P}=[\beta]_{S-P}$ is also a geodesic in $\operatorname{\mathcal{NC}}(S-P)$. Without loss of generality, we may assume that $\nu_{i}$’s and $o$ are in their minimal positions in $S-P$, by taking a hyperbolic metric making $o$ a geodesic in $S-P$ as per Proposition 3.4, which in particular implies that $\nu_{i}$’s are transverse to $o$ and are contained in $\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$. Hence the subgraph distance between $\alpha,\beta^{\prime}\in\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$ is exactly their distance induced from $\operatorname{\mathcal{NC}}^{\dagger}(S)$.
Now combine the results of two paragraphs above, we can conclude that the inclusion map $\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)\hookrightarrow\operatorname{\mathcal{NC}}^{\dagger}(S)$ is a $(1,2)$-quasi-isometry and thus by Proposition 2.3, their Gromov boundaries are homeomorphic. ∎
4 Connectivity properties
In order to demonstrate the connectedness of the Gromov boundary of the fine curve graph, we need only to prove that the spheres in the subgraph $\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$ enjoys connectivity properties as in the assumption of Proposition 2.11.
For notations, for any positive integer $r>0$, let $\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$ be the sphere in $\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$ centred at $o$ of radius $r$, i.e. $\{\alpha\in\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)\colon d^{\dagger}(\alpha,o)=r\}$.
Lemma 4.1.
Let $r>0$ be any positive integer and let $\alpha,\beta\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$. If $\alpha$ and $\beta$ are not transverse, then there exists a $\beta^{\prime}\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$ isotopic to $\beta$ but disjoint from it, which is also transverse to $\alpha$.
Proof.
In light of Proposition 3.5, we can find a finite subset $P\subset S$ so that $(o,\alpha)$ and $(o,\beta)$ are in minimal positions in $S-P$. Up to a small perturbation as in Lemma 3.7, the curve $\beta$ is isotopic to but disjoint from $\beta^{\prime}$ while $\beta^{\prime}$ is transverse to $\alpha$. Since the perturbation can be taken so that $\beta^{\prime}$ remains in a minimal position with $o$ in $S-P$. By Proposition 3.5, $\beta^{\prime}\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$. ∎
Lemma 4.2.
For any positive integer $r>0$, the sphere $\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$ is connected.
Proof.
Let $\alpha,\beta\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$ be distinct non-separating curves. By Lemma 4.1, up to perturbation, we can choose $\beta^{\prime}\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$ isotopic to but disjoint from $\beta$. Moreover, there exists a finite set $P\subset S$ such that $\beta^{\prime}$, $\alpha$ and $o$ are in minimal positions in $S-P$. By taking a hyperbolic metric on $S-P$ as in Proposition 3.4, we can even assume that $\beta^{\prime}$ (up to perturbation), $\alpha$ and $o$ are geodesics. Moreover, in view of Proposition 3.5, there is an isometric embedding via geodesic representations $\iota\colon\operatorname{\mathcal{NC}}(S-P)\hookrightarrow\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$. Note that $S-P$ is of at least genus $2$ and $1$ puncture, i.e. of high complexity as called in [Wri23]. Then applying [Wri23, Proposition 5.4] and [Wri23, Lemma 2.1] to $\operatorname{\mathcal{NC}}(S-P)$, there is a path $[\alpha]_{S-P}=[\nu_{0}]_{S-P},\ [\nu_{1}]_{S-P},\ \dots,\ [\nu_{n}]_{S-P}=[\beta^{\prime}]_{S-P}$ lying inside of the $r$-sphere centred at $[o]_{S-P}$ in $\operatorname{\mathcal{NC}}(S-P)$. Hence the image under the isometric embedding $\iota$ via geodesic representations, $\alpha=\nu_{0},\ \nu_{1},\ \dots,\ \nu_{n}=\beta^{\prime}$, yields a path connecting $\alpha$ to $\beta^{\prime}$ in $\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$. Since $\beta$ and $\beta^{\prime}$ is disjoint, they are adjacent in $\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$. Hence the concatenation $\alpha=\nu_{0},\ \nu_{1},\ \dots,\ \nu_{n}=\beta^{\prime},\ \beta$ is a path in $\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$ connecting the two vertices. ∎
Similar arguments will soon yield Theorem 1.2:
Proof of Theorem 1.2.
For any two distinct $\beta,\gamma\in\{\alpha\in\operatorname{\mathcal{NC}}^{\dagger}(S)\colon d^{\dagger}(\alpha,o)=r\}$, as in Lemma 4.1, we can similarly perturb them to $\beta^{\prime}$ and $\gamma^{\prime}$ that are isotopic to but disjoint from $\beta$ and $\gamma$ respectively, so that $\{o,\beta^{\prime},\gamma^{\prime}\}$ are transverse and $d^{\dagger}(o,\beta^{\prime})=d^{\dagger}(o,\gamma^{\prime})=r$. Now by Lemma 4.2, we can connect them by path contained in the sphere. ∎
The following lemma shows that there is no “dead ends” in the sense of [BM15] inside of the subgraph $\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$, i.e. this subgraph satisfies the assumption (G1) in Proposition 2.11.
Lemma 4.3.
For every adjacent pair $\alpha,\beta\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$ there exists a path
$\alpha=\nu_{0},\nu_{1},\nu_{2}=\beta$
in $\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$ with $\nu_{1}\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r+1)$.
Proof.
Since $\alpha$ and $\beta$ are disjoint, take a non-empty finite subset $P\subset S$ so that $P$ is disjoint from $\alpha\cup\beta\cup o$ and that $o,\alpha,\beta$ are in minimal positions. Take a hyperbolic metric on $S-P$ as in Proposition 3.4 so that $\alpha$ and $\beta$ are geodesics. By [Wri23, Proposition 5.4], we can find a geodesic curve $\nu$ with respect to the above hyperbolic metric such that it is disjoint from $\alpha$ and $\beta$ with $d_{\operatorname{\mathcal{NC}}(S-P)}([\nu]_{S-P},[o]_{S-P})=r+1$. Hence it lies in a connected component $\Sigma$ of $S-\alpha\cup\beta$. Now perturb $\nu$ isotopically into $\nu_{1}$ so that it has no bigon with $o\cap\Sigma$ in $\Sigma-P$. This will make sure that $\nu_{1}$ is still disjoint from $\alpha$ and $\beta$, but is in minimal position with $o$ in $S-P$. Note that $[\nu]_{S-P}=[\nu_{1}]_{S-P}$. Hence $d^{\dagger}(o,\nu_{1})=d_{\operatorname{\mathcal{NC}}(S-P)}([\nu_{1}]_{S-P},[o]_{S-P})=r+1$ by Proposition 3.5. ∎
Lemma 4.4.
Let $r>0$ and $\alpha,\alpha^{\prime}\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$ be isotopic. Then there exists a path on $\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$ connecting $\alpha$ to $\alpha^{\prime}$, in which every vertices is a curve isotopic to $\alpha$. If $\beta\in\operatorname{\mathcal{NC}}^{\dagger}(S)$ is in addition disjoint from both $\alpha$ and $\alpha^{\prime}$, then so will be the vertices in this path.
Proof.
Let $P\subset S$ be a finite subset so that $(o,\alpha)$ and $(o,\alpha^{\prime})$ are in minimal positions in $S-P$. Up to a small perturbation as in Lemma 3.7, we can take $\alpha^{\prime\prime}$ that is isotopic to but disjoint from $\alpha$ and that is transverse to $\alpha^{\prime}$. Since the perturbation is small, the curve $\alpha^{\prime\prime}$ will remain in minimal position with $o$ on $S-P$ and is disjoint from $\beta$. If $\alpha^{\prime\prime}$ is disjoint from $\alpha^{\prime}$, then the proof is done. Otherwise they have bigons. At least one of these bigons bounds a topological disk and contains exactly one subarc of $\alpha^{\prime}$ and one subarc of $\alpha^{\prime\prime}$. By deforming isotopically the subarc of $\alpha^{\prime\prime}$ to the other side of $\alpha^{\prime}$, the curve $\alpha^{\prime\prime}$ can be deformed into an isotopic $\alpha_{1}$ with strictly less intersection points with $\alpha^{\prime}$ and also disjoint from $\alpha^{\prime\prime}$. Since the support of this deformation on $S$ is disjoint from $P$ and $\beta$, so $\alpha_{1}$ remains in minimal position with $o$ and is still disjoint from $\beta$. Hence we can conclude that $d^{\dagger}(\alpha_{1},o)=r$ and $d^{\dagger}(\alpha_{1},\alpha^{\prime\prime})=1$. Continuing the construction process, we gain $\alpha_{1},\ \alpha_{2},\ \dots,\ \alpha_{n}$ until $\alpha_{n}$ is disjoint from $\alpha^{\prime}$ in a way that $d^{\dagger}(\alpha_{i},o)=r$ and $d^{\dagger}(\alpha_{i},\alpha_{i+1})=1$ for all $11\leq i\leq n-$ 1. Hence we get a desired path $\alpha,\alpha^{\prime\prime},\alpha_{1},\dots,\alpha_{n},\alpha^{\prime}$. ∎
If we write $B_{r}(\gamma)\coloneqq\{\alpha\in\operatorname{\mathcal{NC}}^{\dagger}(S)\colon d^{\dagger}(\alpha,\gamma)\leq r\}$ the ball of radius $r$ centred at $\gamma$, then we will have the following result:
Lemma 4.5.
For every $\gamma\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$ and $\alpha,\beta\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r+1)\cap B_{1}(\gamma)$, then there exists a path
$\alpha=\nu_{0},\ \nu_{1},\ \dots,\ \nu_{\ell}=\beta$
in the graph $\operatorname{\mathcal{NC}}^{\dagger}(S)$ with
$\nu_{i}\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r+1)\cap B_{1}(\gamma)$
for $0\leq i\leq\ell$.
Proof.
Let $P\subset S$ be a non-empty finite subset so that $\{o,\gamma,\alpha\}$ are in minimal positions. If $\alpha$ and $\beta$ are isotopic in $S-P$, then by Lemma 4.4, we are done. Suppose now that they are not. Take a hyperbolic metric on $S-P$ so that $o$ and $\gamma$ become geodesic as per Proposition 3.4. Now we take the geodesic representatives $\alpha^{\prime}\in[\alpha]_{S-P}$ and $\beta^{\prime}\in[\beta]_{S-P}$. So $\{o,\gamma,\alpha^{\prime},\beta^{\prime}\}$ are pairwise in minimal position on $S-P$. In particular, we conclude that $\alpha^{\prime},\beta^{\prime}\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r+1)\cap B_{1}(\gamma)$ by Proposition 3.5. By sending the isotopy classes to their geodesic representatives on $S-P$, we can construct an isometric embedding $\iota\colon\operatorname{\mathcal{NC}}(S-P)\hookrightarrow\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$. We remark that [Wri23, Proposition 5.4] implies that there exists a path in $\operatorname{\mathcal{NC}}(S-P)$ connecting $[\alpha^{\prime}]_{S-P}$ to $[\beta^{\prime}]_{S-P}$ in which each vertices has distance $r+1$ to $[o]_{S-P}$ and distance $1$ to $[\gamma]_{S-P}$. Their images under $\iota$ then yield a path from $\alpha^{\prime}$ to $\beta^{\prime}$ included in $B_{1}(\gamma)$. Now we connect a path from $\alpha$ to $\alpha^{\prime}$ as well as from $\beta$ to $\beta^{\prime}$ as in Lemma 4.4. Because the vertices from these two paths are disjoint from $\gamma$, they also lie in $B_{1}(\gamma)$. By consequences, the concatenation of the three paths above yield the desired path. ∎
Now we have finished the proof of all assumptions in Proposition 2.11 for the subgraph $\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)\subset\operatorname{\mathcal{NC}}^{\dagger}(S)\subset\mathcal{C}^{\dagger}(S)$. Note that they are all quasi-isometric. To prove Theorem 1.1, it suffices to apply Proposition 2.11 to $\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$.
Proof of Theorem 1.1.
With $D=1$, for the subgraph $\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$, Lemma 4.3 verifies the assumption (G1) and (G2) from Proposition 2.11 and Lemma 4.5 verifies the assumption (G3) from Proposition 2.11. ∎
Remark 4.6.
If the base point $o$ is chosen as a separating curve on $S$, then instead of linear connectedness, we will obtain a polynomial connectedness, i.e. any two points on the boundary $\xi,\eta\in\partial\mathcal{C}^{\dagger}(S)$ can be included in a connected compact set with diameter at most $L\rho_{o,b}(\xi,\eta)+(L+1)b$ for some $L>0$, since the visual metrics verify the inequality $|\rho_{o,b}-\rho_{o^{\prime},b}|\leq bd^{\dagger}(o,o^{\prime})$.
5 Group action on the boundary
In this section, let us assume that the curves are only $\mathcal{C}^{0}$-embedding of $S^{1}$ into $S$, i.e. we will be considering the topological version of the fine curve graphs. As per Remark 3.2, there is no difference in terms of their boundary at infinity.
Let $S$ be an orientable compact surface as above and let $\operatorname{Homeo}(S)$ be the homeomorphism group of $S$. In this section, we will make several observations on the action of the topological group $\operatorname{Homeo}(S)$ on the topological space $\partial\mathcal{C}^{\dagger}(S)$.
We remark that $\operatorname{Homeo}(S)$ acts by isometries on the fine curve graph $\mathcal{C}^{\dagger}(S)$ and thus on the Gromov bordification of $\mathcal{C}^{\dagger}(S)$. But the action of $\operatorname{Homeo}(S)$ on $\mathcal{C}^{\dagger}(S)$ is not continuous: a sequence of homeomophisms $g_{n}\to\operatorname{Id}$ in compact-open topology (and even isotopic to $\operatorname{Id}$) can send a simple closed curve $\gamma\subset S$ to $g_{n}\gamma$ such that $g_{n}\gamma$ is disjoint from $\gamma$ for all $n$, i.e. $d^{\dagger}(g_{n}\gamma,\gamma)=1$ does not converge to 00 as $n\to\infty$. But since we are only interested in its geometry at large scale, this does not cause any problem:
Proposition 5.1.
The action of $\operatorname{Homeo}(S)$ on $\partial\mathcal{C}^{\dagger}(S)$ is continuous.
Proof.
It suffices to show that whenever given a sequence $g_{n}\to\operatorname{Id}$ as $n\to\infty$, for any $M>0$ and $\xi\in\partial\mathcal{C}^{\dagger}(S)$, there exists $M^{\prime}>0$ and $N>0$ such that if $\langle\xi,\eta\rangle_{o}>M^{\prime}$, then $\langle\xi,g_{n}\eta\rangle_{o}>M$ for any $n>N$. Indeed, since the open sets on the boundary is also generated by shadows defined in (4), we can find an $\alpha\in\mathcal{C}^{\dagger}(S)$ such that whenever $\eta\in\partial\mathcal{C}^{\dagger}(S)$ satisfies $\langle\alpha,\eta\rangle_{o}>R$, we will have $\langle\eta,\xi\rangle_{o}>M$. Moreover, we can assume that $\langle\alpha,\xi\rangle_{o}$ is much larger than $R$. Since the topology on $\operatorname{Homeo}(S)$ is the compact-open topology, there exists $N>0$ such that $g_{n}\alpha$ is contained in a uniformly small neighbourhood of $\alpha$ for every $n>N$. This implies that for any curve $\gamma\subset S$ that is disjoint from this small neighbourhood and a fortiori from $\alpha$, $\gamma$ is also disjoint from $g_{n}\alpha$ for all $n>N$. Hence $d^{\dagger}(\alpha,g_{n}\alpha)\leq 2$ for all $n>N$. The same arguments hold for $o$ so that $d^{\dagger}(o,g_{n}o)\leq 2$ for $n>N$. Let $\eta\in\partial\mathcal{C}^{\dagger}(S)$ such that $\langle\eta,\xi\rangle_{o}>M^{\prime}$. We compute by (2) and (3)
$\langle\alpha,g_{n}\eta\rangle_{o}$ $=\langle g_{n}^{-1}\alpha,\eta\rangle_{g_{n}^{-1}o}\geq\langle g_{n}^{-1}\alpha,\eta\rangle_{o}-2-O(\delta)$
$\geq\min\big{(}\langle\xi,g_{n}^{-1}\alpha\rangle_{o},\langle\xi,\eta\rangle_{o}\big{)}-2-O(\delta)$
$\geq\min\big{(}\langle\xi,g_{n}^{-1}\alpha\rangle_{o},M^{\prime}\big{)}-2-O(\delta)$
But as $d^{\dagger}(\alpha,g_{n}\alpha)\leq 2$, we have $\left|\langle\xi,g_{n}^{-1}\alpha\rangle_{o}-\langle\xi,\alpha\rangle_{o}\right|\leq 2+O(\delta)$, which means that $\langle\xi,g_{n}^{-1}\alpha\rangle_{o}$ is also much larger than $R$. Hence by letting $M^{\prime}$ be large enough, we can assure that $\langle\alpha,g_{n}\eta\rangle_{o}>R$, which implies that $\langle\xi,g_{n}\eta\rangle_{o}>M$ for all $n>N$. ∎
Recall that for a $\delta$-hyperbolic space $X$ and a subgroup $G<\operatorname{Isom}(X)$. Let $o\in X$ be a (and thus any) point in $X$. The limit set of $G$, denoted $\Lambda(G)$, is given by
$\Lambda(G)\coloneqq\left\{\xi\in\partial X\colon g_{n}o\to\xi\ \mathrm{for\ some\ }(g_{n})\in G^{\mathbb{N}}\right\}.$
Now let us examine the limit set of $\operatorname{Homeo}(S)$ on the Gromov boundary of $\mathcal{C}^{\dagger}(S)$.
Lemma 5.2.
Let $S$ be an orientable surface as above. The limit set of $\operatorname{Homeo}(S)$ on $\partial\mathcal{C}^{\dagger}(S)$ is $\partial\mathcal{C}^{\dagger}(S)$.
Proof.
Let us consider $\operatorname{\mathcal{NC}}^{\dagger}(S)$, which is quasi-isometric to $\mathcal{C}^{\dagger}(S)$. We remark that the homeomorphism $\partial\operatorname{\mathcal{NC}}^{\dagger}(S)\to\partial\mathcal{C}^{\dagger}(S)$ is $\operatorname{Homeo}(S)$-equivariant. So it suffices to show that the limit set of $\operatorname{Homeo}(S)$ on $\partial\operatorname{\mathcal{NC}}^{\dagger}(S)$ is $\partial\operatorname{\mathcal{NC}}^{\dagger}(S)$ itself. Indeed, by the theorem of classification of finite-type surfaces, for any two non-separating simple closed curves $\alpha,\beta\subset S$, the bordered surfaces $S-\alpha$ and $S-\beta$ are homeomorphic, so there exists $\phi\in\operatorname{Homeo}(S)$ such that $\phi(\alpha)=\beta$, i.e. the action of $\operatorname{Homeo}(S)$ on $\operatorname{\mathcal{NC}}^{\dagger}(S)$ is transitive. Let $o\in\operatorname{\mathcal{NC}}^{\dagger}(S)$ be any base point and $\xi\in\partial\operatorname{\mathcal{NC}}^{\dagger}(S)$. If $\alpha_{n}$ is a Cauchy-Gromov sequence converging to $\xi$, then by transitivity of the action, for each $n>0$, there exists $\phi_{n}\in\operatorname{Homeo}(S)$ such that $\phi_{n}(o)=\alpha_{n}$. But this means that $\phi_{n}(o)\to\xi$. Hence we can conclude that $\Lambda\big{(}\operatorname{Homeo}(S)\big{)}$ is the entire boundary. ∎
Recall that an isometric group action on a $\delta$-hyperbolic space $X$ is of general type if there exists two independent hyperbolic isometries, or equivalently if the action admits no fixed point on $\partial X$. For isometric actions of general type on $\delta$-hyperbolic spaces, there is an interesting dynamical property:
Proposition 5.3 (Corollary 7.4.3 (ii), [DSU17]).
Let $X$ be a $\delta$-hyperbolic space and let $G$ act on $X$ by isometries. Suppose in addition that the action is of general type. Then the limit set $\Lambda(G)$ is the smallest non-empty closed $G$-invariant subset of $\partial X$.
Recall that a topological space $X$ on which $G$ acts is said $G$-minimal if for every point $x\in X$, the closure of the orbit $\overline{Gx}$ is $X$. Such an action will also be called a minimal action.
Proposition 5.4.
Let $S$ be an orientable surface as above. Then the action of $\operatorname{Homeo}(S)$ on $\partial\mathcal{C}^{\dagger}(S)$ is minimal.
Proof.
In view of Proposition 5.3 and Lemma 5.2, it suffices to show that the action of $\operatorname{Homeo}(S)$ on $\mathcal{C}^{\dagger}(S)$ is of general type. To prove so, it suffices to fix a hyperbolic metric on $S$ and take two independent pseudo-Anosov isometric homeomorphisms on $S$, which will yield two independent hyperbolic isometries on $\mathcal{C}^{\dagger}(S)$ as their action of a subgraph of $\mathcal{C}^{\dagger}(S)$ isometric to $\mathcal{C}(S)$ is of general type (see [MM99]). ∎
Remark 5.5.
The action of $\mathrm{Homeo}(S)$ on $\partial\mathcal{C}^{\dagger}(S)$ is not transitive since the action cannot send the points on $\partial\mathcal{C}^{\dagger}(S)$ that are represented by ending laminations to those who are not.
To conclude this section, we wish to point out in the interests of curiosity that if $S$ is a compact surface, then $\operatorname{Homeo}(S)$ is a Polish group, i.e. a completely metrizable and separable topological group, and its continuous minimal actions on compacta are of special interests. For example, [GTZ21] shows that its universal minimal flow, a compact space with a minimal $\operatorname{Homeo}(S)$-action that admits any compact minimal space for $\operatorname{Homeo}(S)$-actions as a factor, is not metrizable.